# Mode-sum prescription for renormalized expectation values for a charged quantum scalar field on a charged black hole

### v1

@@ Appendix A Coefficients in the Hadamard expansion
In this appendix, we give explicitly the coefficients of the Hadamard parametrix when expanded in “extended coordinates” (with $\Delta r=0$), that is, the coefficients appearing in Eqs. (26) and (26). The assumptions are that the spacetime is a static, spherically symmetric black hole spacetime with surface gravity $\kappa$ and that there is a background electrostatic potential. Otherwise, the metric function $f(r)$ and the Euclideanized potential $A_{\tau}$ are arbitrary functions of $r$. We list the coefficients required to expand the parametrix up to the order such that $\mathcal{O}(\epsilon^{4}\log\,\epsilon)$ terms are ignored, with $\Delta x\sim\epsilon$. Subtracting the parametrix up to this order is sufficient to regularize the stress-energy tensor for a massive charged scalar in an arbitrary static spherically, symmetric spacetime. In practice, we subtract a much higher-order expansion of the parametrix which serves to accelerate the convergence of the mode-sum representation of the renormalized stress-energy tensor. The higher order coefficients are too lengthy to be of use in print form, but we include in the Supplemental Material a Mathematica Notebook with these expressions specialised to Ricci flat spacetimes and imposing the Einstein-Maxwell equations on the vector potential. Below, all derivatives ($f^{\prime}$, $f^{(n)}$ etc.) are with respect to the radius $r$. We omit the explicit dependence on $r$ throughout for typographical convenience.
@@ Appendix B Details of the on-horizon calculations
$q_{0l}(r)=q^{\textrm{approx}}_{0l}(r)+\beta_{l}p_{0l}(r)$ (96)
@@
$q^{\textrm{approx}}_{0l}(r)=\frac{\Upsilon^{1/4}}{[r^{2}f(r)]^{1/4}}K_{0}(\rho_{l}\sqrt{\Upsilon})$

### v2

@@ Appendix A Coefficients in the Hadamard expansion
In this appendix, we give explicitly the coefficients of the Hadamard parametrix when expanded in “extended coordinates” (with $\Delta r=0$), that is, the coefficients appearing in Eqs. (27) and (27). The assumptions are that the space-time is a static, spherically symmetric black hole space-time with surface gravity $\kappa$ and that there is a background electrostatic potential. Otherwise, the metric function $f(r)$ and the Euclideanized potential $A_{\tau}$ are arbitrary functions of $r$. We list the coefficients required to expand the parametrix up to the order such that $\mathcal{O}(\epsilon^{4}\log\,\epsilon)$ terms are ignored, with $\Delta x\sim\epsilon$. Subtracting the parametrix up to this order is sufficient to regularize the stress-energy tensor for a massive charged scalar in an arbitrary static spherically, symmetric space-time. In practice, we subtract a much higher-order expansion of the parametrix which serves to accelerate the convergence of the mode-sum representation of the renormalized stress-energy tensor. The higher order coefficients are too lengthy to be of use in print form, but we include in the Supplemental Material a Mathematica Notebook with these expressions specialised to Ricci flat space-times and imposing the Einstein-Maxwell equations on the vector potential. Below, all derivatives ($f^{\prime}$, $f^{(n)}$ etc.) are with respect to the radius $r$. We omit the explicit dependence on $r$ throughout for typographical convenience.
@@ Appendix B Details of the on-horizon calculations
$q_{0l}(r)=q^{\textrm{approx}}_{0l}(r)+\beta_{l}p_{0l}(r),$ (98)
@@
$q^{\textrm{approx}}_{0l}(r)=\frac{\Upsilon^{1/4}}{[r^{2}f(r)]^{1/4}}K_{0}(\rho_{l}\sqrt{\Upsilon}),$ (99)